# Any-point Trajectory Modeling for Policy Learning

### 16 Feb 2024 v2

@@ Abstract
Figure 1: Given a task instruction and the initial positions of any set of points in an image frame, our Any-point Trajectory Model (ATM) can predict the future trajectories of these points conditioned on the task. After training the model on a video dataset, the predicted trajectories serve as effective guidance for learning visuomotor policies for a set of language-conditioned manipulation tasks.
@@ I Introduction
Computer vision and natural language understanding have recently made significant advances [20, 7], largely due to the availability of large datasets. Similarly, in robotics, scaling up human demonstration data has been key for learning new skills [6, 32, 13], with a clear trend of performance improvement with larger datasets [27, 6]. However, human demonstrations, typically action-labeled trajectories collected via teleoperation devices [52, 49], are time-consuming and labor-intensive to collect. For instance, collecting $130K$ trajectories in [6] took 17 months, making data collection a major bottleneck in robot learning.
@@
Videos contain knowledge about behaviors, physics, and semantics, presenting an alternative data source. However, the lack of action labels makes utilization of video data in policy learning difficult. Previous works have addressed this by using self-supervised objectives for video pre-training to learn a feature representation of the observation for policy learning [41, 31, 39]. However, a feature representation only describes the state at the current time step, largely neglecting the transition dynamics that predicts future states. To explicitly model the transition dynamics, prior works have developed video prediction models that predict future image frames from current ones [11, 51, 12]. However, learning a video prediction model for control introduces two issues. Firstly, the task of video prediction avoids any abstraction by modeling changes to every pixel, coupling the physical motion with visual appearances such as texture, and lighting. This coupling makes modeling difficult, often resulting in hallucinations and unrealistic future predictions [11]. Secondly, these models are computationally demanding in both training and inferencing. With limited computational resources, performance significantly declines. Moreover, the high inference cost compels these models to adopt open-loop execution [11, 5], which tends to result in less robust policies.
@@
In this paper, we propose a novel and structured representation to bridge video pre-training and policy learning. We first represent each state as a set of points in a video frame. To model the temporal structure in videos, we learn an Any-point Trajectory Model (ATM) that takes the positions of the points in the current frame as input and outputs their future trajectories. We predict these trajectories in the camera coordinate frame to minimize the assumptions on calibrated cameras. These 2D point trajectories correspond to trajectories of particles in the 3D space and are a universal representation of the motions that can transfer to different domains and tasks. Contrasting with the video prediction approach of tracking changes in pixel intensity, our particle-based trajectory modeling offers a more faithful abstraction of the physical dynamics, naturally incorporating inductive biases like object permanence. We first pre-train the trajectory model on actionless video datasets. After pre-training, the predicted trajectories serve as detailed guidance for the policies, functioning like subgoals. We then train trajectory-guided policies using only a minimal amount of action-labeled demonstration data. For training the ATMs, we generate self-supervised training data by leveraging recent advancements in vision models for accurate point tracking [18]. Across over 130 language-conditioned tasks we evaluated in both simulation and the real world, ATM significantly surpasses various strong baselines in video pre-training, achieving an average success rate of $63\%$ compared to the highest success rate of $37\%$ by previous methods, marking an improvement of over $80\%$. Additionally, we demonstrate that our method facilitates effective transfer learning from human videos and videos of a robot with a different morphology. We summarize our main contributions below:
@@
    We demonstrate effective learning from human videos.
@@ II Related Work
Video pre-training for control. Videos contain rich information about behaviors and dynamics, which can help policy learning. However, video pre-training remains challenging due to the lack of action labels. One line of works first learns an inverse dynamics model that predicts the action from two adjacent frames and then labels the videos with pseudo actions [37, 3, 44]. However, the inverse dynamics model is often trained on a limited action-labeled dataset and does not generalize well, especially for continuous actions. Prior works have also explored pre-train a feature representation using various self-supervised objectives [40, 31, 26], but the representation alone does not retain the temporal information in videos. More recently, learning video prediction as pre-training has shown promising results [11, 21, 4, 51, 12]. During policy learning, a video prediction model is used to generate future subgoals and then a goal-conditioned policy can be learned to reach the sub-goal. However, video prediction models often result in hallucinations and unrealistic physical motions. Furthermore, video models require extensive computation, which is an issue, especially during inference time. In contrast, our method models the trajectories of the points, naturally incorporating inductive bias like object permanence while requiring much less computation. This enables our trajectory models to be run closed-loop during policy execution. Furthermore, the trajectories provide dense guidance to the policy as a motion prior.
Learning from Human Videos. Of particular interest, numerous prior works have proposed learning from the rich source of human videos [50, 42, 30, 2, 43]. However, these works often explicitly extract the hand pose or contact regions from the human videos, thereby losing information about the dynamics of the remaining objects. In contrast, our method models the trajectories of arbitrary points and can learn from both human videos and videos of a different robot.
@@ III Preliminary
Tracking Any Point (TAP). The recent advancements in video tracking [9, 10, 47] enable us to track the trajectory of each point in video frames without external supervision. In this paper, we utilize the off-the-shelf tracker proposed in [18]. Formally, given a sequence of images from a video $o_{1},...,o_{T}$, any one of the time steps $\bar{t}\in[1,T]$, and a set of points in that frame $\{p_{\bar{t},k}\}_{k=1}^{K}$, where $p_{\bar{t},k}=(x,y)$ is the point coordinate in the camera frame, the task of tracking is to predict the 2D camera-frame coordinates of the corresponding points in every frame $p_{t,k}$ where $t=1\dots T$. In this paper, we use the terms trajectory and track interchangeably to refer to a sequence of 2D coordinates of any point, denoted as $(p_{1},\dots,p_{T})$. We only model the 2D trajectories in the camera frame so that we do not have to make additional assumptions about multi-view cameras, or the availability of depth, allowing future scaling to more diverse video datasets. The tracker additionally predicts a binary visibility value $v_{t,k}$ denoting whether the point is occluded at step $t$.
@@ IV Method
We propose Any-point Trajectory Modeling (ATM). As illustrated in Figure 2, ATM learns to predict future point trajectories in a video frame as the pre-training, then uses the predicted trajectories to guide policy learning. Our proposed method will be comprehensively detailed in this section: In Sec. IV-A, we first describe how to learn a point trajectory prediction model from an action-free video dataset $\mathcal{T}_{o}$. Then in Sec. IV-B, we outline how we utilize the pre-trained track prediction model to learn a track-guided policy from a limited action-labeled trajectory datasets $\mathcal{T}_{a}$.
@@ IV-B Track-guided Policy Learning
A track-guided policy $\pi(a_{t}|o_{t},\textbf{p}_{t:t+H})$ takes input the current observation $o_{t}$ and the predicted tracks $\textbf{p}_{t:t+H}$ and predict the actions. A simplified illustration of our policy architecture is shown in Figure 3. Our transformer policy architecture follows the architecture in prior works [25, 19]. Although the predicted tracks alone already provide rich information to predict the actions, we still incorporate contextual image observations into our policy so that no information is lost, as suggested in prior works [24]. We incorporate the track tokens both before and after fusion with the image tokens (early fusion and late fusion) to ensure that the guiding information from tracks can be effectively integrated. Surprisingly, as the tracks already provide the fine-grained subgoals, we find that the policy no longer needs language instruction at this stage as task specification. Essentially, the provided tracks have reduced the difficult policy learning problems into a much easier sub-goal following problems, reducing the policy into an inverse dynamics model. Our track-guided policy is trained with MSE loss. A detailed architecture diagram and hyperparameters are available in the appendix.
@@ V Experiments
  • How does ATM compare with state-of-the-art video pre-training baselines in learning from action-free videos?
@@ V-A Video Pre-training for Imitation Learning
    BC denotes the vanilla behavioral cloning which trains a policy exclusively using the limited action-labeled expert demonstrations, without using the video dataset. It uses a policy architecture identical to ATM except that the particle trajectories are masked to be zero and it instead takes language embedding as task specification.
@@
Results. We present the main results in Figure 4. We see that by bridging the video data and policy learning with the structured representation of point trajectories, ATM (our method) significantly surpasses various strong baselines in video pre-training, achieving an average success rate of $63\%$ compared to the highest success rate of $37\%$ by previous methods, marking an improvement of over $80\%$. The comparison of BC with ATM shows that learning from additional videos provides useful information for policy learning. VPT performs poorly as we empirically observed that the pseudo-action labels predicted by VPT generally show large errors on the video dataset. UniPi fails on more complex as video prediction models are not physically grounded and often generate future frames that are not physically feasible, such as cases where robots disappear from the image. Please see our video for failure cases of a video prediction model.
@@ V-C Ablation Analysis
We conduct a series of ablation experiments in simulation to demonstrate the effect of our design choices. These include the number of action-labeled trajectories needed for policy learning, the effect of the trajectory prediction horizon, image masking, and late track fusion.
@@
[…]
Figure 9: We plot the success rates of the policies learned with predicted trajectories of different lengths. Generally, longer trajectory length improves the performance, but the benefit tends to plateau at 16.
@@
Effect of image masking in track transformer. When training the track transformer, we randomly mask out patches in the images and learn to reconstruct them as an auxiliary task. We conduct an ablation study by removing the image masking loss and comparing it against our standard configuration, where we randomly mask out $50\%$ of the image masks. In Table II, the results reveal a slight decline in policy performance when image masking is omitted, suggesting image masking can be a useful auxiliary task.
@@ A-A Simulation Experiments
Attention map visualization. To demonstrate how tracks guide the policy, we visualize the attention maps between the spatial CLS token and RGB tokens in the spatial transformer of BC and our method. Figure 10 demonstrates that our method effectively focuses on the relevant spatial regions, as specified by the textual instructions. Specifically, it attends to the cream cheese, bowl, and wine bottle in the respective example tasks, while BC is usually distracted by the irrelevant regions, highlighting the superior capability of the tracks as better task prompts.
@@ A-B Discussions on the UniPi Baselines
TABLE IV: Computation cost for performing a single task consisting of 600 time steps for different methods. ATM performs trajectory generation at each time step while being computationally efficient. UniPi performs an open-loop future goal generation once at the beginning using a video diffusion model. UniPi-Replan generates image sub-goals at more frequent time steps and is the slowest.
Method ATM UniPi UniPi-Replan
Computation (TFLOPS) 1.56 39.29 2946.75
@@
A-C Human-to-robot Transfer Details
@@
TABLE VI: Hyperparameters of track transformer training.
TABLE VII: Hyperparameters of policy training.
@@
TABLE VII: Hyperparameters of policy training.

### 12 Jul 2024 v3

@@ Abstract
Figure 1: Given a task instruction and the initial positions of any set of points in an image frame, our Any-point Trajectory Model (ATM) can predict the future trajectories of these points conditioned on the task. After training the model on an action-free video dataset, the predicted trajectories serve as effective guidance for learning visuomotor policies for a set of language-conditioned manipulation tasks.
@@ I Introduction
Computer vision and natural language understanding have made significant advances in recent years [22, 7], where the availability of large datasets plays a critical role. Similarly, in robotics, scaling up human demonstration data has been key for learning new skills [6, 34, 14], with a clear trend of performance improvement with larger datasets [29, 6]. However, human demonstrations, typically action-labeled trajectories collected via teleoperation devices [56, 53], are time-consuming and labor-intensive to collect. For instance, collecting $130K$ trajectories in [6] took 17 months, making data collection a major bottleneck in robot learning.
@@
Videos contain knowledge about behaviors, physics, and semantics, presenting an alternative data source. However, the lack of action labels makes utilization of video data in policy learning difficult. Previous works have addressed this by using self-supervised objectives for video pre-training to learn a feature representation of the observation for policy learning [44, 33, 42]. However, a feature representation only describes the state at the current time step, largely neglecting the transition dynamics that predicts future states. To explicitly model the transition dynamics, prior works have developed video prediction models that predict future image frames from current ones to guide policy learning [12, 55, 13]. However, learning a video prediction model for control introduces two challenges. Firstly, the task of video prediction avoids any abstraction by modeling changes to every pixel, coupling the physical motion with visual appearances such as texture, and lighting. This coupling makes modeling difficult, often resulting in hallucinations and unrealistic future predictions [12]. Secondly, these models are computationally demanding in both training and inferencing. With limited computational resources, performance significantly declines. Moreover, the high inference cost compels these models to adopt open-loop execution [12, 5], which tends to result in less robust policies.
@@
In this paper, we propose a novel and structured representation to bridge video pre-training and policy learning. We first represent each state as a set of points in a video frame. To model the temporal structure in videos, we learn an Any-point Trajectory Model (ATM) that takes the positions of the points in the current frame as input and outputs their future trajectories. We predict these trajectories in the camera coordinate frame to minimize the assumptions on calibrated cameras. These 2D point trajectories correspond to trajectories of particles in the 3D space and are a universal representation of the motions that can transfer to different domains and tasks. Contrasting with the video prediction approach of tracking changes in pixel intensity, the particle-based trajectory modeling offers a faithful abstraction of the physical dynamics, naturally incorporating inductive biases like object permanence and continuous motion. We first pre-train the trajectory model on actionless video datasets. After pre-training, the predicted trajectories can function as subgoals to guide the control policy. We then train trajectory-guided policies using only a minimal amount of action-labeled demonstration data. For training the ATMs, we generate self-supervised training data by leveraging recent advancements in vision models for accurate point tracking [20]. Across over 130 language-conditioned tasks we evaluated in both simulation and the real world, ATM significantly surpasses various strong baselines in video pre-training, achieving an average success rate of $63\%$ compared to the highest success rate of $37\%$ by previous methods, marking an improvement of over $80\%$. Additionally, we demonstrate that our method facilitates effective transfer learning from human videos and videos of a robot with a different morphology. We summarize our main contributions below:
@@
    We demonstrate effective learning from cross-embodiment human and robot videos.
@@ II Related Work
Video pre-training for control. Videos contain rich information about behaviors and dynamics, which can help policy learning. However, video pre-training remains challenging due to the lack of action labels. One line of works first learns an inverse dynamics model that predicts the action from two adjacent frames and then labels the videos with pseudo actions [39, 3, 47, 40]. However, the inverse dynamics model is often trained on a limited action-labeled dataset and does not generalize well, especially for continuous actions. Prior works have also explored pre-train a feature representation using various self-supervised objectives [43, 33, 28], but the representation alone does not retain the temporal information in videos. More recently, learning video prediction as pre-training has shown promising results [12, 23, 4, 55, 13]. During policy learning, a video prediction model is used to generate future subgoals and then a goal-conditioned policy can be learned to reach the sub-goal. However, video prediction models often result in hallucinations and unrealistic physical motions. Furthermore, video models require extensive computation, which is an issue, especially during inference time. In contrast, our method models the trajectories of the points, naturally incorporating inductive bias like object permanence while requiring much less computation. This enables our trajectory models to be run closed-loop during policy execution. Furthermore, the trajectories provide dense guidance to the policy as a motion prior.
Learning from Human Videos. Of particular interest, numerous prior works have proposed learning from the rich source of human videos [54, 45, 32, 2, 46, 50]. However, these works often explicitly extract the hand pose or contact regions from the human videos, thereby losing information about the dynamics of the remaining objects. In contrast, our method models the trajectories of arbitrary points and can learn from both human videos and videos of a different robot.
@@ III Preliminary
Tracking Any Point (TAP). The recent advancements in video tracking [17, 10, 11, 51, 57] enable us to track the trajectory of each point in video frames without external supervision. In this paper, we utilize the off-the-shelf tracker proposed in [20]. Formally, given a sequence of images from a video $o_{1},...,o_{T}$, any one of the time steps $\bar{t}\in[1,T]$, and a set of points in that frame $\{p_{\bar{t},k}\}_{k=1}^{K}$, where $p_{\bar{t},k}=(x,y)$ is the point coordinate in the camera frame, the task of tracking is to predict the 2D camera-frame coordinates of the corresponding points in every frame $p_{t,k}$ where $t=1\dots T$. In this paper, we use the terms trajectory and track interchangeably to refer to a sequence of 2D coordinates of any point, denoted as $(p_{1},\dots,p_{T})$. We only model the 2D trajectories in the camera frame so that we do not have to make additional assumptions about multi-view cameras, or the availability of depth, allowing future scaling to more diverse video datasets. The tracker additionally predicts a binary visibility value $v_{t,k}$ denoting whether the point is occluded at step $t$.
@@ IV Method
We propose Any-point Trajectory Modeling (ATM). As illustrated in Figure 2, ATM is a two-stage framework: first learn to predict future point trajectories in a video frame as the pre-training with large-scale action-free videos, then use the predicted trajectories to guide policy learning with few-shot action-labeled demos. Our proposed method will be comprehensively detailed in this section: In Sec. IV-A, we first describe how to learn a point trajectory prediction model from an action-free video dataset $\mathcal{T}_{o}$. Then in Sec. IV-B, we outline how we utilize the pre-trained track prediction model to learn a track-guided policy from a limited action-labeled trajectory datasets $\mathcal{T}_{a}$.
@@ IV-B Track-guided Policy Learning
A track-guided policy $\pi(a_{t}|o_{t},\textbf{p}_{t:t+H})$ takes input the current observation $o_{t}$ and the predicted tracks $\textbf{p}_{t:t+H}$ and predict the actions. A simplified illustration of our policy architecture is shown in Figure 3. Our transformer policy architecture follows the architecture in prior works [27, 21]. Although the predicted tracks alone already provide rich information to predict the actions, we still incorporate contextual image observations into our policy so that no information is lost, as suggested in prior works [26]. We incorporate the track tokens both before and after fusion with the image tokens (early fusion and late fusion) to ensure that the guiding information from tracks can be effectively integrated. Surprisingly, as the tracks already provide the fine-grained subgoals, we find that the policy no longer needs language instruction at this stage as task specification. Essentially, the provided tracks have reduced the difficult policy learning problems into a much easier sub-goal following problems, reducing the policy into an inverse dynamics model. Our track-guided policy is trained with MSE loss. Note that, the weights of our policy model are randomly initialized rather than copied from the pretrained Track Transformer like other video-pretraining methods [33, 28], because of the aim to separately study the task guidance capability of predicted trajectories. A detailed architecture diagram and hyperparameters are available in the appendix.
@@ V Experiments
  • How does ATM compare with state-of-the-art video pre-training and behaviour cloning baselines for learning from action-free videos?
@@ V-A Video Pre-training for Imitation Learning
    BC denotes the vanilla behavioral cloning which trains a policy exclusively using the limited action-labeled expert demonstrations, without using the video dataset. It uses a policy architecture identical to ATM except that the particle trajectories are masked to be zero and it instead takes language embedding as task specification. BC baseline can also be considered as an ablation for policy learning without action-free videos. The specific architecture and losses for training the BC policy can vary. We use BC to denote a transformer policy trained using MSE loss by default and compare with diffusion policy separately towards the end of this section.
@@
Results. We present the main results in Figure 4. We see that by bridging the video data and policy learning with the structured representation of point trajectories, ATM (our method) significantly surpasses various strong baselines in video pre-training, achieving an average success rate of $63\%$ compared to the highest success rate of $37\%$ by previous methods, marking an improvement of over $80\%$. The comparison of BC with ATM shows that learning from additional videos provides useful information for policy learning. VPT performs poorly as we empirically observed that the pseudo-action labels predicted by VPT generally show large errors on the video dataset. UniPi fails on more complex as video prediction models are not physically grounded and often generate future frames that are not physically feasible, such as cases where robots disappear from the image. ATM shows surprisingly superior performance on long-horizon tasks and tasks that require an understanding of goals. We attribute this performance gain to ATM’s formulation, which utilizes explicit future particle trajectories as subgoals. At each time step, ATM predicts a future trajectory goal based on the current observation. In comparison to the static natural language instructions used by other methods, which do not change throughout the episode, ATM’s closed-loop future trajectory provides clear guidance as to what the policy needs to do next. Please see our video for failure cases of a video prediction model.
Universality of ATM on Different Policy Classes. The benefit of the predicted trajectories can also be shown for state-of-the-art policy class such as diffusion policy [8], as shown in Figure 6. Here, ATM Diffusion Policy adds the predicted future tracks as additional condition to the policy class, while keeping everything else the same. We see that ATM consistently improves the performance of diffusion policy across all the suites, achieving the highest scores on the LIBERO benchmark. More details can be found in Appendix.
Figure 6: We implement ATM Diffusion Policy by adding the predicted future trajectories as additional conditioning and show consistent improvement over the base diffusion policies across the benchmark suites.
@@ V-C Ablation Analysis
We conduct a series of ablation experiments in simulation to demonstrate the effect of our design choices. These include the number of action-labeled trajectories needed for policy learning, the effect of the trajectory prediction horizon, image masking, and late track fusion. Besides, we also investigate the universality of ATM for different policy models.
@@
Effect of the number of action-labeled trajectories. We investigate the impact of track length on our policy input by training our policy using predicted tracks of varying lengths: 4, 8, 16, 32, and 64 steps. As illustrated in Figure 10, a track length of 16 yields optimal performance on average. Smaller track lengths significantly decrease performance, while a track length of 0 reduces ATM to the behavior cloning baseline without any benefit of pre-training. On the other hand, larger track lengths can also negatively impact performance, especially on more challenging tasks such as LIBERO-Long and LIBERO-Goal, as longer track predictions are noisier and less relevant for achieving short-horizon subgoals. Determining the optimal time horizon for subgoals could be an interesting avenue for future research.
Figure 10: We plot the success rates of the policies learned with predicted trajectories of different lengths. Generally, longer trajectory length improves the performance, but the benefit tends to plateau after 16.
@@
Effect of image masking in track transformer. When training the track transformer, we randomly mask out patches in the images and learn to reconstruct them as an auxiliary task. We conduct an ablation study by removing the image masking loss and comparing it against our standard configuration, where we randomly mask out $50\%$ of the image masks. In Table II, the results reveal a slight decline in policy performance when image masking is omitted, suggesting image masking can be a useful auxiliary task, except on LIBERO-Spatial. We hypothesize that the auxiliary loss encourages the model to jointly reason about different regions of an image observation. This bias may not be critical for the Libero-spatial tasks because the specific spatial location of interest is already fully specified in the language instruction (e.g., all tasks have the format of ”pick up the black bowl at [some spatial location]”).
@@ A-A Simulation Experiments
Attention map visualization of track-guided policy. To demonstrate how tracks guide the policy, we visualize the attention maps between the spatial CLS token and RGB tokens in the spatial transformer of BC and our method. Figure 11 demonstrates that our method effectively focuses on the relevant spatial regions, as specified by the textual instructions. Specifically, it attends to the cream cheese, bowl, and wine bottle in the respective example tasks, while BC is usually distracted by the irrelevant regions, highlighting the superior capability of the tracks as better task prompts.
@@ A-B Discussions on the UniPi Baselines
TABLE IV: Computation cost and inference time for different methods on a V100 GPU. ATM performs trajectory generation instead of predicting high-dimensional frames, making it the most computationally efficient and feasible for closed-loop control. UniPi employs a video diffusion model for open-loop future goal generation at the beginning, demanding significantly higher computational resources. UniPi-Replan generates fewer frames than UniPi using a smaller model, resulting in marginally faster generation. However, its use in closed-loop control remains computationally prohibitive.
Computation Close-Loop Open-Loop
ATM UniPi-Replan UniPi
TFLOPS per generation 1.56 13.09 39.29
Time per generation (s) 0.015 4.51 8.14
@@
A-C Applying ATM to Diffusion Policy
In addition to the transformer policy adopted from the original LIBERO paper [27], our ATM can also be integrated with the state-of-the-art Diffusion Policy [8]. Specifically, we implement an ATM Diffusion Policy by using predicted tracks as input conditions. Both the standard and ATM-based Diffusion policies are trained with 20% of the data from LIBERO, in the same setting as the experiment in Figure 4.
As shown in Table VI, the standard Diffusion Policy achieves impressive results on the LIBERO benchmark, even with limited training data, confirming its status as the state-of-the-art imitation policy. Moreover, the ATM Diffusion Policy consistently outperforms the standard model, highlighting our framework’s versatility and effectiveness across different policy architectures.
TABLE VI: Detailed results of Diffusion policy on LIBERO. The Diffusion policy can be further improved by our method, suggesting that our Any-point Trajectory Modeling framework is an important building block to apply to any policy model.
Diffusion Policy $67.67\pm 1.25$ $78.00\pm 2.45$ $35.00\pm 3.74$ $37.33\pm 2.05$ $33.85\pm 1.71$
ATM Diffusion Policy $79.00\pm 3.74$ $81.00\pm 2.45$ $58.67\pm 4.64$ $44.00\pm 6.38$ $62.89\pm 1.10$
A-D Human-to-robot Transfer Details
@@
Furthermore, we visualize the attention maps of Track Transformers trained with or without human videos in Figure 14. Track Transformer trained with human videos (bottom row) displays well-defined attention heatmaps that primarily focus on the object and robot arm. Conversely, the attention maps from the model trained without human videos (middle row) are notably more dispersed, often incorrectly focusing on background areas. This comparison highlights the effectiveness of our ATM in leveraging prior knowledge from human videos, facilitating successful cross-embodiment imitation learning.
TABLE VII: Hyperparameters of track transformer training.
@@
Figure 14: Attention maps of Track Transformer trained with and without human videos. Including large-scale human video leads to much clearer attention maps, focusing on the object and robot arm, while training without human videos will attend to incorrect areas such as background walls.